I cannot understand how to do a problem like this. It is simplifying powers.
(-2x)(-3xy^2)^2. I am not sure what I multiply together.
I know the answer is -18x^3y^4. I am trying to review stuff I had done a while back, but cannot figure this out. Thanks in advance for any help and understanding you can offer.

Answer by sudhanshu_kmr(1152)

(-2x)(-3xy^2)^2
= (-2x)(-3)^2*x^2*(y^2)^2
= (-2x)*9*x^2*y^4
= -2*9*x*x^2*y^4
= -18x^3y^4

always remember......
for (-a)^n if n is even then: a^n
if n is odd then: -a^n

a^x * a^y = a^(x+y)

(a^x)^y = a^(x*y)